A spring becomes 15 cm long when the force of 40 N is used to pull the spring. The elastic potential energy stored in the spring is 2.2 J. What is the original length of the spring?


1
Expert's answer
2021-08-27T14:40:13-0400

Given:

"l=0.15\:\rm m"

"F=40\:\rm N"

"E_p=2.2\:\rm J"


The lastic potential energy stored in the spring 

"E_p=\frac{kx^2}{2}=\frac{Fx}{2}"

So, the elongation of a spring

"x=l-l_0=\frac{2E_p}{F}=\frac{2*2.2}{40}=0.11\:\rm m"

Finally, the original length of the spring

"l_0=l-x=0.15-0.11=0.04\:\rm m=4\:\rm cm"